OpenStudy (anonymous):

what value of y makes the equation true (-9.7*24) * (-0.25)=-9.7 * (24*y)

OpenStudy (anonymous):

i don't now this one

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

this question can be solved by using associative property of the multiplication \[(a*b)*c = a*(b*c)\]

OpenStudy (anonymous):

still dont get it

OpenStudy (anonymous):

\[(a*b)*c =a*(b*c)\] \[(-9.7*24)*(-0.25) =-9.7*(24*y)\] so whats the value of y